# MMhops-R1: Multimodal Multi-hop Reasoning

### v1

@@ Introduction
With continuous advancement in reasoning capabilities, Large Language Models (LLMs) like OpenAI’s o1 (Jaech et al. 2024), DeepSeek-R1 (Guo et al. 2025), and Kimi-k1.5 (Team et al. 2025) demonstrate strong performance in complex problem-solving by extending chain-of-thought reasoning during inference. Multimodal large language models (MLLMs), by inheriting the reasoning abilities or adopting similar training paradigms, achieve significant progress in integrating visual understanding and language reasoning (Xu et al. 2024; Peng et al. 2025; Zhang et al. 2025; Zheng et al. 2025). However, current multimodal reasoning research primarily focuses on stimulating intrinsic model capabilities, such as spatial reasoning (Zhou et al. 2025), object detection (Chen et al. 2025a; Liu et al. 2025), and mathematical reasoning (Meng et al. 2025; Leng et al. 2025). By contrast, complex real-world problems typically require the integration of multimodal reasoning with external knowledge retrieval through multi-turn interactions, enabling multimodal multi-hop reasoning. For example, Figure 1(B1) requires the model to extract information from the image, retrieve relevant external knowledge, and perform multi-step reasoning to reach the answer. Figure 1(B2) requires identifying entities across multiple images, retrieving corresponding knowledge, and conducting quantitative reasoning.
Despite progress, existing Visual Question Answering (VQA) datasets remain insufficient for multimodal multi-hop reasoning. Current datasets are limited in both visual and textual reasoning depth: standard VQA datasets typically require only single-step visual understanding (Goyal et al. 2017; Hudson and Manning 2019; Singh et al. 2019), while several knowledge-based VQA datasets (Marino et al. 2019; Schwenk et al. 2022; Lerner et al. 2022; Chen et al. 2023) introduce external knowledge retrieval to increase complexity. However, as shown in Figure 1(A), models usually use one step of visual recognition and one step of text retrieval to answer, without constructing complex reasoning chains. E-VQA (Mensink et al. 2023) extends questions to two-hop reasoning, but this extension remains restricted to the textual domain and features a fixed reasoning path length, lacking multimodal integration and diverse reasoning types. These limitations make existing datasets inadequate for effectively supporting model training and evaluation in complex multimodal multi-hop reasoning tasks.
@@
Knowledge-Based VQA. To advance VQA beyond perception towards more complex reasoning, the task of KB-VQA was introduced, which requires models to incorporate external knowledge. However, prominent KB-VQA datasets like OK-VQA (Marino et al. 2019) and A-OKVQA (Schwenk et al. 2022) were largely confined to commonsense knowledge or simple facts. Subsequent efforts, including ViQuAE (Lerner et al. 2022) and INFOSEEK (Chen et al. 2023), expanded the knowledge domain to large-scale corpora such as Wikipedia. Nonetheless, these datasets predominantly feature questions solvable via a two-step process: identifying a visual entity and executing a single query against a knowledge base. While the recent E-VQA (Mensink et al. 2023) dataset introduced textual two-hop reasoning, its reasoning chains are confined to the textual modality and a fixed length. In contrast, the MMhops dataset is the first to systematically require multi-hop reasoning across both visual and textual modalities, featuring diverse, variable-length reasoning paths, demanding a more profound integration of multi-modal information.
Multimodal and Multi-hop RAG. Early mRAG frameworks (Caffagni et al. 2024; Yan and Xie 2024; Zhang et al. 2024) typically employ a static, single-step pipeline: they first retrieve relevant documents based on the initial query and then feed them to the generator. A key limitation of these approaches is their reliance on a static, pre-defined process, which lacks the flexibility to adapt to queries of varying complexity. While recent work like OmniSearch (Comanici et al. 2025) introduces a planning agent, it relies on manually engineered prompts or supervised fine-tuning, which does not equip the model with the intrinsic capability to learn complex reasoning policies autonomously. In parallel, multi-hop RAG has emerged in the unimodal text domain to address similar challenges. To move beyond fixed reasoning chains, methods such as Search-R1 (Jin et al. 2025) and ReSearch (Chen et al. 2025b) leverage reinforcement learning (RL) with algorithms like GRPO (Shao et al. 2024) and PPO (Schulman et al. 2017) to train an agent that learns a dynamic retrieval policy. However, these powerful RL-based paradigms have thus far been confined to the textual modality. Our work, MMhops-R1, bridges this divide by extending this RL-based paradigm to the mRAG domain, training an agent to strategically orchestrate retrieval and reasoning across both visual and textual knowledge sources.
@@
OK-VQA (Marino et al. 2019) 14K 1 1 2 Factoid
A-OKVQA (Schwenk et al. 2022) 24.9K 1 1 2 Common sense/Factoid
ViQuAE (Lerner et al. 2022) 3.7K 1 1 2 Wikipedia
INFOSEEK (Chen et al. 2023) 1.35M 1 1 2 Wikipedia
E-VQA (Mensink et al. 2023) 1M 1 1-2 2-3 Wikipedia
@@
The MMhops dataset is built based on the Wikipedia Knowledge Base (KB), with an automated data annotation and quality filtering process designed using powerful language models like GPT-4o (Hurst et al. 2024). The dataset includes two core reasoning types: Bridging Questions and Comparison Questions, covering various reasoning depths and different numbers of image inputs.
@@
    Entity Collection: Collect a large number of visual entities from the Wikipedia knowledge base. Use the embedding model NV-Embed-v2 (Lee et al. 2024) to perform semantic similarity matching based on the entity names and summary, selecting entities with high relevance. Use LLMs to perform semantic deduplication and remove pairs of entities that refer to the same concept.
@@
GPT-4o-mini (Hurst et al. 2024) 29.67 20.54 26.08 23.80 7.05
GPT-4o (Hurst et al. 2024) 41.66 33.60 39.28 36.62 8.76
Gemini-2.5-flash (Comanici et al. 2025) 51.08 43.51 50.10 46.58 23.18
Gemini-2.5-pro (Comanici et al. 2025) 58.80 50.83 57.32 53.98 29.39
@@
Search-r1 (Jin et al. 2025) Qwen2.5-7b-Instruct Caption, Text 14.45 23.05 17.85 19.98 6.62
Self-Ask (Press et al. 2022) GPT-4o Caption, Text 27.59 31.41 31.13 30.42 18.27
@@
EchoSight (Yan and Xie 2024) LLaMA3 Image, Text 19.14 11.83 11.86 13.63 4.81
OmniSearch (Li et al. 2024) GPT-4o Image, Text 31.02 49.77 36.5 42.65 17.02
@@
To optimize our policy $\pi_{\theta}$ using a composite reward signal, we adapt the objective function from DAPO (Yu et al. 2025). This objective is coupled with a dynamic sampling strategy that filters generated response groups. Specifically, we enforce that each group of $G$ responses must contain at least one factually correct sample, as stipulated by the constraint in our objective. This design enables the policy to optimize for procedural correctness by creating non-zero advantages ($\hat{A}_{i,t}$) from process-based rewards, such as format adherence ($R_{\text{format}}$), even when the final outcome is already correct. Our full optimization objective is formulated as:
@@
We optimize our policy using the Verl framework (Sheng et al. 2024), employing Qwen2.5-VL-7B-Instruct (Bai et al. 2025) as the backbone model. The model is trained for a single epoch on the MMhops dataset with a constant learning rate of $1\times 10^{-6}$. During policy optimization, we use a batch size of 256 and a group size of 8. Our knowledge base for retrieval comprises 100K Wikipedia articles, each accompanied by an image. For image retrieval, we utilize the CLIP-ViT-L/14@336px model (Radford et al. 2021). For text retrieval, we employ the E5 model (Wang et al. 2022) to fetch the top-3 most relevant passages for each query. The maximum number of interaction turns with the knowledge base is set to 4 during both training and inference. The hyperparameters $\alpha$, $\beta$, and $\gamma$ for the reward function are set to 1.0, 1.0, and 0.25, respectively.
@@
We adopt the evaluation protocol from INFOSEEK (Chen et al. 2023), categorizing answers into three types: String, Time, and Numerical. For String answers, we report Exact Match (EM) accuracy. For Time answers, we employ EM with a tolerance of $\pm 1$ year. For Numerical answers, a prediction is deemed correct if it falls within a $\pm 0.1$ margin of the ground truth or achieves an Intersection-over-Union (IoU) of at least 50%. The overall score is the weighted average of the accuracies for each type. We report performance on the Test set, with a breakdown for bridging and comparison questions.
@@
CLIP-PaLM (Chen et al. 2023) 22.7 18.5 20.4
CLIP-FiD (Chen et al. 2023) 23.3 19.1 20.9
Wiki-LLaVA (Caffagni et al. 2024) 30.1 27.8 28.9
EchoSight (Yan and Xie 2024) 31.3
@@
To verify the generalizability of the proposed method, we evaluate it on two widely used knowledge-based VQA datasets: INFOSEEK (Chen et al. 2023) and E-VQA (Mensink et al. 2023). Results on INFOSEEK show the effectiveness of MMhops-R1 on multimodal single-hop questions, while its performance on two-hop questions in E-VQA confirms its generalization ability to multi-hop reasoning.

### v2

@@ Introduction
With continuous advancement in reasoning capabilities, Large Language Models (LLMs) like OpenAI’s o1 (jaech2024openai), DeepSeek-R1 (guo2025deepseek), and Kimi-k1.5 (team2025kimi) demonstrate strong performance in complex problem-solving by extending chain-of-thought reasoning during inference. Multimodal large language models (MLLMs), by inheriting the reasoning abilities or adopting similar training paradigms, achieve significant progress in integrating visual understanding and language reasoning (xu2024llava; peng2025skywork; zhang2025r1; zheng2025deepeyes). However, current multimodal reasoning research primarily focuses on stimulating intrinsic model capabilities, such as spatial reasoning (zhou2025r1), object detection (chen2025vinci; liu2025visual), and mathematical reasoning (meng2025mm; leng2025mmr1). By contrast, complex real-world problems typically require the integration of multimodal reasoning with external knowledge retrieval through multi-turn interactions, enabling multimodal multi-hop reasoning. For example, Figure 1(B1) requires the model to extract information from the image, retrieve relevant external knowledge, and perform multi-step reasoning to reach the answer. Figure 1(B2) requires identifying entities across multiple images, retrieving corresponding knowledge, and conducting quantitative reasoning.
Despite progress, existing Visual Question Answering (VQA) datasets remain insufficient for multimodal multi-hop reasoning. Current datasets are limited in both visual and textual reasoning depth: standard VQA datasets typically require only single-step visual understanding (goyal2017making; hudson2019gqa; singh2019towards), while several knowledge-based VQA datasets (marino2019ok; schwenk2022okvqa; lerner2022viquae; chen2023can) introduce external knowledge retrieval to increase complexity. However, as shown in Figure 1(A), models usually use one step of visual recognition and one step of text retrieval to answer, without constructing complex reasoning chains. E-VQA (mensink2023encyclopedic) extends questions to two-hop reasoning, but this extension remains restricted to the textual domain and features a fixed reasoning path length, lacking multimodal integration and diverse reasoning types. These limitations make existing datasets inadequate for effectively supporting model training and evaluation in complex multimodal multi-hop reasoning tasks.
@@
Knowledge-Based VQA. To advance VQA beyond perception towards more complex reasoning, the task of KB-VQA was introduced, which requires models to incorporate external knowledge. However, prominent KB-VQA datasets like OK-VQA (marino2019ok) and A-OKVQA (schwenk2022okvqa) were largely confined to commonsense knowledge or simple facts. Subsequent efforts, including ViQuAE (lerner2022viquae) and INFOSEEK (chen2023can), expanded the knowledge domain to large-scale corpora such as Wikipedia. Nonetheless, these datasets predominantly feature questions solvable via a two-step process: identifying a visual entity and executing a single query against a knowledge base. While the recent E-VQA (mensink2023encyclopedic) dataset introduced textual two-hop reasoning, its reasoning chains are confined to the textual modality and a fixed length. In contrast, the MMhops dataset is the first to systematically require multi-hop reasoning across both visual and textual modalities, featuring diverse, variable-length reasoning paths, demanding a more profound integration of multi-modal information.
Multimodal and Multi-hop RAG. Early mRAG frameworks (caffagni2024wiki; yan2024echosight; zhang2024mr) typically employ a static, single-step pipeline: they first retrieve relevant documents based on the initial query and then feed them to the generator. A key limitation of these approaches is their reliance on a static, pre-defined process, which lacks the flexibility to adapt to queries of varying complexity. While recent work like OmniSearch (comanici2025gemini) introduces a planning agent, it relies on manually engineered prompts or supervised fine-tuning, which does not equip the model with the intrinsic capability to learn complex reasoning policies autonomously. In parallel, multi-hop RAG has emerged in the unimodal text domain to address similar challenges. To move beyond fixed reasoning chains, methods such as Search-R1 (jin2025search) and ReSearch (chen2025learning) leverage reinforcement learning (RL) with algorithms like GRPO (shao2024deepseekmath) and PPO (schulman2017proximal) to train an agent that learns a dynamic retrieval policy. However, these powerful RL-based paradigms have thus far been confined to the textual modality. Our work, MMhops-R1, bridges this divide by extending this RL-based paradigm to the mRAG domain, training an agent to strategically orchestrate retrieval and reasoning across both visual and textual knowledge sources.
@@
OK-VQA (marino2019ok) 14K 1 1 2 Factoid
A-OKVQA (schwenk2022okvqa) 24.9K 1 1 2 Common sense/Factoid
ViQuAE (lerner2022viquae) 3.7K 1 1 2 Wikipedia
INFOSEEK (chen2023can) 1.35M 1 1 2 Wikipedia
E-VQA (mensink2023encyclopedic) 1M 1 1-2 2-3 Wikipedia
@@
The MMhops dataset is built based on the Wikipedia Knowledge Base (KB), with an automated data annotation and quality filtering process designed using powerful language models like GPT-4o (hurst2024gpt). The dataset includes two core reasoning types: Bridging Questions and Comparison Questions, covering various reasoning depths and different numbers of image inputs.
@@
    Entity Collection: Collect a large number of visual entities from the Wikipedia knowledge base. Use the embedding model NV-Embed-v2 (lee2024nv) to perform semantic similarity matching based on the entity names and summary, selecting entities with high relevance. Use LLMs to perform semantic deduplication and remove pairs of entities that refer to the same concept.
@@
GPT-4o-mini (hurst2024gpt) 29.67 20.54 26.08 23.80 7.05
GPT-4o (hurst2024gpt) 41.66 33.60 39.28 36.62 8.76
Gemini-2.5-flash (comanici2025gemini) 51.08 43.51 50.10 46.58 23.18
Gemini-2.5-pro (comanici2025gemini) 58.80 50.83 57.32 53.98 29.39
@@
Search-r1 (jin2025search) Qwen2.5-7b-Instruct Caption, Text 14.45 23.05 17.85 19.98 6.62
Self-Ask (press2022measuring) GPT-4o Caption, Text 27.59 31.41 31.13 30.42 18.27
@@
EchoSight (yan2024echosight) LLaMA3 Image, Text 19.14 11.83 11.86 13.63 4.81
OmniSearch (li2024benchmarking) GPT-4o Image, Text 31.02 49.77 36.5 42.65 17.02
@@
To optimize our policy $\pi_{\theta}$ using a composite reward signal, we adapt the objective function from DAPO (yu2025dapo). This objective is coupled with a dynamic sampling strategy that filters generated response groups. Specifically, we enforce that each group of $G$ responses must contain at least one factually correct sample, as stipulated by the constraint in our objective. This design enables the policy to optimize for procedural correctness by creating non-zero advantages ($\hat{A}_{i,t}$) from process-based rewards, such as format adherence ($R_{\text{format}}$), even when the final outcome is already correct. Our full optimization objective is formulated as:
@@
We optimize our policy using the Verl framework (sheng2024hybridflow), employing Qwen2.5-VL-7B-Instruct (bai2025qwen2) as the backbone model. The model is trained for a single epoch on the MMhops dataset with a constant learning rate of $1\times 10^{-6}$. During policy optimization, we use a batch size of 256 and a group size of 8. Our knowledge base for retrieval comprises 100K Wikipedia articles, each accompanied by an image. For image retrieval, we utilize the CLIP-ViT-L/14@336px model (radford2021learning). For text retrieval, we employ the E5 model (wang2022text) to fetch the top-3 most relevant passages for each query. The maximum number of interaction turns with the knowledge base is set to 4 during both training and inference. The hyperparameters $\alpha$, $\beta$, and $\gamma$ for the reward function are set to 1.0, 1.0, and 0.25, respectively.
@@
We adopt the evaluation protocol from INFOSEEK (chen2023can), categorizing answers into three types: String, Time, and Numerical. For String answers, we report Exact Match (EM) accuracy. For Time answers, we employ EM with a tolerance of $\pm 1$ year. For Numerical answers, a prediction is deemed correct if it falls within a $\pm 0.1$ margin of the ground truth or achieves an Intersection-over-Union (IoU) of at least 50%. The overall score is the weighted average of the accuracies for each type. We report performance on the Test set, with a breakdown for bridging and comparison questions.
@@
CLIP-PaLM (chen2023can) 22.7 18.5 20.4
CLIP-FiD (chen2023can) 23.3 19.1 20.9
Wiki-LLaVA (caffagni2024wiki) 30.1 27.8 28.9
EchoSight (yan2024echosight) 31.3
@@
To verify the generalizability of the proposed method, we evaluate it on two widely used knowledge-based VQA datasets: INFOSEEK (chen2023can) and E-VQA (mensink2023encyclopedic). Results on INFOSEEK show the effectiveness of MMhops-R1 on multimodal single-hop questions, while its performance on two-hop questions in E-VQA confirms its generalization ability to multi-hop reasoning.